# Local map Construction Methods with SD map: A Novel Survey

### v1

@@ 3 Multimodal Representation
3.1 Image Data
@@
In the perception task of BEV, the image information of the panoramic camera is the most important input data, and the common feature extraction method of the panoramic image follows the paradigm of autonomous driving perception task BEVformer[20] or LSS[21]. The Backbone module of neural networks extracts 2D image features from various camera perspectives through classic and lightweight convolutional networks such as ResNet-50[22] or 101, Mobilenets[23], EfficientNet[24], V2-99[25] and so on. Among them, the ResNet series is widely used because of ResNet architecture addresses the vanishing gradient problem in deep neural networks during training by introducing residual blocks. Variants like ResNet enhance feature extraction capabilities by increasing network depth and width. These networks are extensively utilized in BEV Local map perception tasks due to their outstanding performance in image recognition and feature extraction. Typically, a Feature Pyramid Network (FPN)[26] module is appended to the Backbone module. The FPN integrates feature maps of different scales, generating more robust multi-scale feature representations. This seems to be the default basic configuration, and the number of fusion layers can be selected according to the network type. This multi-scale feature fusion aids in improving the detection and recognition of objects of varying sizes, thereby enhancing overall performance.
@@
In addition to these lightweight and simple backbone, larger backbone networks will be the mainstream trend in the future. With the success of Transformers in the field of computer vision[27], Transformer-based feature extraction methods have also been applied to BEV Local map perception tasks such as Swin[28]. Refer to the methods on the Nuscece leaderboard, The state-of-the-art methods all use pre-trained VIT-L as the backbone network, or its variant EVA-02[29]. This large pre-trained backbone network is the key to improving model performance, although the large number of parameters and computational complexity of large models can seriously affect inference speed. Nevertheless, its performance directly promotes detection accuracy. The training of these large models requires massive data support, but data labeling is costly and limited, and self supervised training methods will become mainstream. With the widespread application of BERT[30] pre-trained models in various downstream tasks for self supervised tasks in natural language processing, it has demonstrated a powerful ability to learn language representations. Similarly, in self supervised learning in computer vision tasks, MAE[31] randomly masks patches on images and implements self supervised learning of masked images. The achievements of MIM[32] based pre-training algorithms are flourishing in the field of computer vision. This self supervised pre-training model can not only solve the problem of high cost labels, but also better learn the representation relationship of images.
@@
3.2 Lidar Points Data
In the Local map perception task of BEV, in addition to using a pure visual surround camera as a single data input, multimodal methods fuse multi-modal information such as lidar point clouds and camera data to perform depth aware BEV transformation. Compared to single vision and multimodal (RGB+LiDAR) methods, multimodal fusion methods perform excellently in accuracy despite increasing additional computational complexity. The processing of lidar point cloud data is a crucial step in multimodal perception tasks. The feature extraction of lidar point cloud data in P-mapnet[33] requires voxelization of the point cloud first, followed by the use of multi-layer perceptron (MLP) to extract local features of each point. Maximizing pooling selects the largest feature value from multiple local features to form a global feature representation, enhancing the model’s global perception ability of point cloud data[34].
@@ 3.2 Lidar Points Data
where $\mathit{F_{1}}$ represents the feature extractor, extracting multimodal inputs to obtain BEV features, $\mathit{F_{2}}$ represents the decoder, and outputs the detection results.
The method in MapLite 2.0[35] further integrates lidar point cloud data with data from other sensors and integrates it with coarse road maps obtained from SD maps (such as OpenStreetMap) Use the coarse route map information from the SD map to refine the geometric shape and topological structure of the road. This not only improves the accuracy of the map, but also enhances the understanding of complex road environments [35]. It is also used to generate high-definition maps online by projecting LiDAR intensity data in bird’s-eye view. By integrating multimodal data, not only does it provide detailed spatial information, but it also enables precise semantic segmentation of the driving environment.
3.3 SD Map Data
In the context of enhancing Local map perception tasks, incorporating SD map information as prior knowledge can significantly improve the performance of vision and lidar sensors, particularly in long-distance and occlusion scenarios [2]. To integrate SD maps effectively into network structures while preserving their unique road information, various representations have been explored. SD maps can generally be categorized into two forms: raster and vector.
@@ 3.3 SD Map Data
Figure 2: Sample graph of SD map from nuscenes dataset
Figure 3: Overview of the chronological order of local map construction
@@ 3.3.1 Representation of Raster
MapLite2.0 [9] was the first to introduce SD map into Local map perception tasks. PriorLane [36] modeling the map as a binary image, where 1 represents the drivable area and 0 represents the non drivable area. Similarly, MapVision[37] also uses the one hot encoding method, then concatenates the position encoding information and extracts the SD map features through the encoder. The SD map is aligned with ego data through the KEA module proposed in the article, and then fused with sensor data to obtain a mixed expression. Both P-MapNet and MapLite2.0 use rasterization to represent SD map, but the difference is that after P-MapNet[33], a CNN network is used to extract information from the rasterized SD map, which is used as a source of additional information for BEV feature optimization (i.e. key and val); MapLite2.0 takes SD map as the initial estimate of HDMap, converts it into a BEV perspective and combines it with images input from sensors. It is trained through a Convolutional Neural Network to predict its semantic labels. Finally, these semantic segmentation results are transformed into distance transformations for specific labels, and a structured estimator is used to maintain Local map estimates and integrate SD map priors.
@@ 3.3.2 Representation of Vectors
SMERF[2] was the first to propose a Transformer based encoder model for road topology inference. MapEX[38] and SMERF have similar representations of map elements, introducing a ployline sequence representation and a Transformer encoder to obtain the final map representation of the scene. Specifically, the roads in SD map are first abstracted in the form of polylines. For the ployline data, N data points are obtained through uniform sampling. Then, after sin cos encoding, the N-d dimensional line description is obtained. Consider a vertical line with small curvature, which is characterized by very similar x or y axis values for all points. Directly inputting the coordinates of these points into the model may result in insufficient differentiation of this curvature.
@@ 3.3.3 Encoding of Other Information
SMERF: In addition to encoding the polyline coordinates of SD map. SMERF use one hot encoding to encode the type of road into a vector with dimension K (the number of road types). For ground elements within the perceptual range, M * (N * d + K) encoded data will be obtained, which will be transformed through several layers to obtain map feature representations. The ablation experiment has shown that adding more road type information can improve the effectiveness of lane detection and road topology inference.
@@ 4 Multimodal Fusion Method
The method of using only images as input, exemplified by MapTR [39], based on the encoder-decoder architecture, has established a classic paradigm for local map construction, paving the way for subsequent approaches. Streammapnet [7] further enhances this by incorporating comprehensive temporal information, significantly improving performance in occluded areas. 3D LaneNet [40] adopts an end-to-end learning framework, integrating tasks such as image encoding, spatial transformation between image views and top-down views, and 3D curve extraction into a single network. Gen LaneNet [41] proposes a two-stage framework that decouples the learning of image segmentation sub-networks and geometric encoding sub-networks.
@@
Additionally, several monocular 3D lane detection methods, such as [42], [43], and [44], focus solely on visual images as input. Numerous models, including [45], [6], [5], [46], [47], [48], [49], [50], [51], [52], [53], [54], [55], and [56], also rely solely on visual images. On the other hand, HDMapNet [8], a representative multimodal method, integrates LiDAR point clouds by encoding these features and predicting vectorized map elements in bird’s-eye views, achieving effective fusion of multi-sensor data. Furthermore, other models, such as [57], [58], [59], [60], [61], [62], [63], and [64], incorporate LiDAR point cloud data as additional input. Fig. 3 illustrates the development trends in local map construction over recent years. Considering the cost of constructing high-precision maps, Maplite 2.0 [35] was the first to introduce SD maps into local map perception tasks. MapEX [38] addresses situations where existing map information is incomplete or inaccurate by converting existing map elements into non-learnable queries and combining them with learnable queries for training and prediction. SMERF [2] and P-MapNet [33] combine the feature representation of SD maps with camera input features using a multi-head cross-attention mechanism, enabling more effective lane topology inference.
To achieve effective fusion of visual BEV features and SD map semantic information, BLOS-BEV [65] has explored various feature fusion methods. Additionally, methods such as PriorLane [36], FlexMap [66], Bayesian [67], TopoLogic [68], LGMap [69], MapVision [37], RoadPainter [70], and EORN [71] integrate SD map priors to support local map construction, a trend that is gradually gaining traction.
Figure 4: The Overview of MapTR
@@
A perspective transition is required before fusion. The focus of this section is to convert feature information extracted from 2D camera sensor images (commonly referred to as perspective views (PV)) into BEV features. Local map perception tasks typically view the ground as a plane, in the Bird’s Eye View, Establishing a map in BEV, because on one hand, BEV facilitates information fusion of multiple sensors, and existing advanced BEV object detection work can provide a good foundation. There are both geometric based and network-based methods for converting the perspective from PV to BEV. Geometric methods can be divided into those based on isomorphic transformations and those based on depth estimation. Network based methods can be divided into MLP based and Transformer based methods. The conversion from PV to BEV based on Transformer can usually be directly achieved using the BEV perception model. MapTR[39] in Fig. 4 proposes an optimized GTK module based on the View Transformer module in BEVFormer.
4.1 Align

### v2

@@ 3 Multimodal Representation
In this section, our paper will explore the data modalities, including the extraction and processing of image data, point cloud data, and SDMap data, as well as an introduction to the fusion methods for the aforementioned three types of modal data.
3.1 Image
@@
In the perception task of BEV, the image information of the panoramic camera is the most important input data, and the common feature extraction method of the panoramic image follows the paradigm of intelligent driving perception task BEVformer[18] or Lift-Splat-Shoot (LSS)[19]. The backbone module of neural networks extracts 2D image features from various camera perspectives through classic and lightweight convolutional networks such as ResNet-50 or 101[20], MobileNets[21], EfficientNet[22], V2-99[23] and so on. Among them, the ResNet series[20] is widely used and variants like ResNet enhance feature extraction capabilities by increasing network depth and width. These networks are extensively utilized in BEV local map perception tasks due to their outstanding performance in image recognition and feature extraction. Typically, a Feature Pyramid Network (FPN)[24] module is appended to the backbone module. The FPN[24] integrates feature maps of different scales, generating more robust multi-scale feature representations. This seems to be the default basic configuration, and the number of fusion layers can be selected according to the network type. This multi-scale feature fusion aids in improving the detection and recognition of objects of varying sizes, thereby enhancing overall performance.
@@
In addition to lightweight and simple backbone, larger backbone networks will be the mainstream trend in the future. With the success of transformers in the field of computer vision[25], transformer-based feature extraction methods have also been applied to BEV local map perception tasks such as Swin Transformer[26]. Refer to the methods on the nuScenes leaderboard, The state-of-the-art methods all use pre-trained VIT-L[25] as the backbone network, or its variant EVA-02[27]. This large pre-trained backbone network is the key to improving model performance, although the large number of parameters and computational complexity of large models can seriously affect inference speed. Nevertheless, its performance directly promotes detection accuracy.
The training of these large models requires massive data support, but data labeling is costly and limited, and self-supervised training methods will become mainstream. With the widespread application of BERT[28] pre-trained models in various downstream tasks for self-supervised tasks in natural language processing, it has demonstrated a powerful ability to learn language representations. Similarly, in self-supervised learning in computer vision tasks, MAE[29] randomly masks patches on images and implements self-supervised learning of masked images. The achievements of MIM[30] based pre-training algorithms are flourishing in the field of computer vision. This self-supervised pre-training model can not only solve the problem of high-cost labels, but also better learn the representation relationship of images.
@@
3.2 Lidar Points
In the local map perception task of BEV, in addition to using a pure visual surround camera as a single data input, multimodal methods fuse multi-modal information such as lidar point clouds and camera data to perform depth-aware BEV transformation. Compared to single vision and multimodal (RGB+LiDAR) methods, the multimodal fusion method performs excellently in accuracy despite increasing additional computational complexity. The processing of lidar point cloud data is a crucial step in multimodal perception tasks. The feature extraction of lidar point cloud data in P-MapNet[31] requires voxelization of the point cloud first, followed by the use of multi-layer perceptron (MLP) to extract local features of each point. Maximizing pooling selects the largest feature value from multiple local features to form a global feature representation, enhancing the model’s global perception ability of point cloud data[32].
@@ 3.2 Lidar Points
where $\mathit{F_{1}}$ represents the feature extractor, extracting multimodal inputs to obtain BEV features, $\mathit{F_{2}}$ represents the decoder and outputs the detection results.
The method in MapLite 2.0[33] further integrates lidar point cloud data with data from other sensors and integrates it with coarse road map obtained from SDMap (such as OpenStreetMap[8]) Use the coarse route map information from the SDMap to refine the geometric shape and topological structure of the road. This not only improves the accuracy of the map, but also enhances the understanding of complex road environments[33]. It is also used to generate high-definition maps online by projecting lidar intensity data in bird’s-eye view. Integrating multimodal data, not only provides detailed spatial information, but also enables precise semantic segmentation of the driving environment.
3.3 SDMap
In the context of enhancing local map perception tasks, incorporating SDMap information as prior knowledge can significantly improve the performance of vision and lidar sensors, particularly in long-distance and occlusion scenarios[2]. To integrate SDMap effectively into network structures while preserving their unique road information, various representations have been explored. SDMap can generally be categorized into two forms: raster and vector.
@@ 3.3 SDMap
Figure 3: Sample graph of SDMap from nuScenes dataset.
Figure 4: This diagram chronologically presents some of the classic local map construction methods using SDMap in recent years, as well as other methods.
@@ 3.3.1 Representation of Raster
MapLite2.0[34] was the first to introduce SDMap into local map perception tasks. PriorLane[35] modeling the map as a binary image, where ’1’ represents the drivable area and ’0’ represents the non-drivable area. Similarly, MapVision[36] also uses the one-hot encoding method, then concatenates the position encoding information and extracts the SDMap features through the encoder. The SDMap is aligned with ego data through the KEA module proposed in the algorithm and then fused with sensor data to obtain a mixed expression. Both P-MapNet[31] and MapLite2.0[33] use rasterization to represent SDMap, but the difference is that after P-MapNet[31], a CNN is used to extract information from the rasterized SDMap, which is used as a source of additional information for BEV feature optimization (i.e. key and value); MapLite2.0[33] takes SDMap as the initial estimate of HDMap, converts it into a BEV perspective and combines it with images input from sensors. It is trained through a CNN to predict its semantic labels. Finally, these semantic segmentation results are transformed into distance transformations for specific labels, and a structured estimator is used to maintain local map estimates and integrate SDMap priors.
@@ 3.3.2 Representation of Vectors
SMERF[2] was the first to propose a transformer-based encoder model for road topology inference and introduce a ployline sequence representation and a transformer encoder to obtain the final map representation of the scene. Most subsequent studies followed this approach. Specifically, the roads in SDMap are first abstracted in the form of polylines. For the ployline data, N data points are obtained through uniform sampling. Then, after sin cos encoding, the N-d dimensional line description is obtained. Consider a vertical line with small curvature, which is characterized by very similar x-axis or y-axis values for all points. Directly inputting the coordinates of these points into the model may result in insufficient differentiation of this curvature.
@@ 3.3.3 Encoding of Other Information
In addition to encoding the polyline coordinates of SDMap. SMERF[2] uses one-hot encoding to encode the type of road into a vector with dimension K (the number of road types). For ground elements within the perceptual range, M * (N * d + K) encoded data will be obtained, which will be transformed through several layers to obtain map feature representations. The ablation experiment has shown that adding more road type information can improve the effectiveness of lane detection and road topology inference.
@@ 4 Multimodal Fusion Method
The method of using only images as input, exemplified by MapTR[37], based on the encoder-decoder architecture, has established a classic paradigm for local map construction, paving the way for subsequent approaches. MapEX[38] addresses situations where existing map information is incomplete or inaccurate by converting existing map elements into non-learnable queries and combining them with learnable queries for training and prediction. StreamMapNet[7] further enhances this by incorporating comprehensive temporal information, significantly improving performance in occluded areas. 3D-LaneNet[39] adopts an end-to-end learning framework, integrating tasks such as image encoding, spatial transformation between image views and top-down views, and 3D curve extraction into a single network. Gen-LaneNet[40] proposes a two-stage framework that decouples the learning of image segmentation sub-networks and geometric encoding sub-networks.
@@
Additionally, several monocular 3D lane detection methods, such as[41], [42], and[43], focus solely on visual images as input. Numerous models, including[6], [5], [44], [45], [46], [47], [48], [49], [50], [51], [52], [53], [54], [49], [55] and [56], also rely solely on visual images. On the other hand, HDMapNet[57], a representative multimodal method, integrates point clouds by encoding these features and predicting vectorized map elements in BEV, achieving effective fusion of multi-sensor data. Furthermore, other models, such as[58], [59], [60], [61], [62], [63], and [64], incorporate point cloud data as additional input. Fig. 4 illustrates the development trends in local map construction over recent years. Considering the cost of constructing high-precision maps, Maplite 2.0[33] was the first to introduce SDMap into local map perception tasks. SMERF[2] and P-MapNet[31] combine the feature representation of SDMap with camera input features using a multi-head cross-attention mechanism, enabling more effective lane topology inference. To achieve an effective fusion of visual BEV features and SDMap semantic information, BLOS-BEV[65] has explored various feature fusion methods. Additionally, methods such as PriorLane[35], FlexMap[66], Bayesian[67], TopoLogic[68], LGMap[69], MapVision[36], LSM (Leveraging SD Map to Assist the OpenLane Topology)[70], UniHDMap[71], RoadPainter[72], and EORN[73] integrate SDMap priors to support local map construction, a trend that is gradually gaining traction (Table 2).
Table 2: Performance of using SDmap method on dataset OpenLane-V2
Method SDmap representation $DET_{l}$uparrow $DET_{t}$uparrow $DET_{a}$uparrow $TOP_{ll}$uparrow $TOP_{lt}$uparrow $OLS$uparrow $OLUS$uparrow
TopoNet+SMERF[2] vector 33.4 48.6 - 7.5 23.4 39.4 -
TopoLogic[68] - 34.4 48.3 - 23.4 24.4 45.1 -
MapVision[36] vector 39 80 40 38 48 - 58
LGmap[69] vector 50.74 - 55.57 46.32 53.59 - 66
RoadPainter[72] vector 36.9 47.1 - 12.7 25.8 42.6 -
TopoNet+OSMR[73] raster 30.6 44.6 - 7.7 22.9 37.7 -
TopoNet+OSMG[73] vector 30 47.6 - 5.4 21.3 36.7 -
LSM[70] vector 49.97 - 49.80 - - - 63.9
UniHDMap[71] vector 49.94 79.27 46.38 43.92 52.11 - 62.81
Figure 5: The pipeline of MapTR using self-attention variants and cross-attention variants.
@@
A perspective transition is required before fusion. The focus of this section is to convert feature information extracted from 2D camera sensor images (commonly referred to as PV) into BEV features. Local map perception tasks typically view the ground as a plane, in the Bird’s Eye View, Establishing a map in BEV, because on one hand, BEV facilitates information fusion of multiple sensors, and existing advanced BEV object detection work can provide a good foundation. There are both geometric-based and network-based methods for converting the perspective from PV to BEV. Geometric methods can be divided into those based on isomorphic transformations and those based on depth estimation. Network-based methods can be divided into MLP-based and transformer-based methods. The conversion from PV to BEV based on the transformer method can usually be directly achieved using the BEV perception model. MapTR[37] in Fig. 5 uses an optimized GKT[74] module based on the View Transformer module in BEVFormer[18].
PriorLane[35] primarily consists of three successive components: the MiT block, the KEA module, and the FT block. According to the structure depicted in Fig. 6, Unlike typical ViTs[25], the MiT block can generate hierarchical features due to its design of overlapping patch merging and efficient self-attention mechanisms. The Knowledge Embedding Alignment (KEA) module is essential for spatially aligning the knowledge embeddings with the coarse vehicle positions in the image features. The knowledge encoder layer and the fusion encoder layer work together to form the fusion transformer. The FT block introduces a novel approach to detection transformer pre-training by leveraging implicit vision prompts, which in PriorLane[35] are generated from explicit prior knowledge, and notably, no annotated labels are used for prompt training.
Figure 6: The overview of PriorLane, using prior feature of SDMap to fuse with image feature.
PriorLane[35] introduces an innovative strategy to enhance lane detection by combining image features with cost-effective local prior knowledge. This prior knowledge is represented as a BEV grid map, segmented into patches, and converted into learnable embeddings through a trainable linear projection. To address the challenge of spatial alignment, PriorLane[35] incorporates a KEA module, which ensures that the embeddings are accurately aligned spatially, thus preserving consistency between the BEV and camera perspectives. The key innovation of PriorLane[35] lies in its fusion process, employing a transformer-based architecture to integrate the image features, extracted via an encoder, with the prior knowledge embeddings. Inspired by Segformer[75], PriorLane[35] utilizes a MLP block to effectively merge the fused features with the original image features, yielding pixel-level segmentation predictions. This methodology has been validated through comparative experiments, demonstrating significant enhancements in lane detection performance by effectively leveraging prior knowledge in conjunction with image features.
MapLite 2.0[33] converts a coarse road graph into an HDmap prior and a rasterized BEV representation. The system generates additional BEV images from sensor inputs, which are processed using the CNN trained on existing HDMap data to predict semantic labels. These semantic segmentation results are then converted into distance transforms specific to each label. A structured estimator simultaneously updates the local map and integrates the SDMap prior. This approach enables the online generation of HDMap, relying solely on SDMap and onboard sensor data. As a result, autonomous vehicles can navigate unmapped areas, while the system also detects environmental changes that could signal the need for map updates.
SDMap provides lightweight priors, requiring minimal storage and bandwidth. While offering less detail than HDMap, SDMap is sufficient for navigation, representing roads as graphs where nodes indicate intersections and edges represent road centerlines, annotated with details like road names, speed limits, and lane counts. In contrast, the HDmap model is layered into road networks, sections with constant lane counts, and directional lanes, ensuring each section connects to specific nodes.
Figure 7: The overview of MapLite 2.0. Using SDMap as prior information, it supports the online inference of local high-definition maps from perception data.
To create BEV features, MapLite[34] demonstrates that point clouds from lidar sensors are gathered over time and paired with camera images (Fig. 7). These points are then projected into a BEV frame by dividing the space around the vehicle into a grid that captures intensity, height, and RGB values for each grid cell. A ResNet-101-based semantic segmentation model is utilized to predict the positions of important map objects such as lane markings and road boundaries. This model is fine-tuned via transfer learning from a DeepLabv3 network[76] pre-trained on the COCO dataset[77]. After segmentation, the output is converted into signed distance transforms, which handle occlusions and remove artifacts caused by missing data.
In the online HDmap estimation process, initial map parameters are modeled as Gaussian distributions. The Prior of MapLite[34], the first HDmap estimate, is constructed offline by comparing SD and HD maps, predicting parameters based on observed distributions. As new sensor measurements arrive, the map state is updated by incorporating the distance transforms from the segmentation step. Optimization is then used to refine parameter estimates, ensuring that sensor data is fully integrated while maintaining an accurate and reliable map estimate.
The SMERF framework[2] enhances an existing lane topology model by integrating priors from SDMap, leading to improved detection of lane centerlines and enhanced relational reasoning. In this context, the SDMap is encoded into a feature representation through a transformer encoder. By applying cross-attention between the feature representation of the SDMap and the onboard camera inputs, the framework constructs BEV features crucial for lane detection and relational reasoning. This pipeline undergoes end-to-end training with the lane topology model, requiring no additional training signals. To leverage the polyline sequence representation of the SDMap, a transformer encoder is employed to develop a feature representation tailored for the downstream lane topology task. The embedding of the polyline sequence occurs through a linear layer, consistent with typical transformer encoder architectures. This process ensures that the discrete one-hot representation of road types is effectively transformed into a continuous space.
Figure 8: The overview of SMERF. Its SDMap feature is encoded from the polylines and road types of SDMap.
As shown in Fig. 8, SMERF[2] fuses the features derived from the SDMap with intermediate BEV feature representations by utilizing multi-head cross-attention. This method is adaptable to nearly all transformer-based lane topology models, as it establishes cross-attention between the BEV feature queries and SDMap features at each intermediate encoder layer. The fused BEV features not only incorporate 3D information obtained from images but also capture road-level geometric details extracted from the SDMap. Subsequently, the decoder of the lane topology model takes these SDMap-augmented features as input to predict lane centerlines, traffic elements, and affinity matrices that facilitate the association between lane centerlines and traffic elements. Importantly, the integration of SDMap empowers the model to infer lanes that are distant or even obscured by buildings.
In P-MapNet[31], the localized SDMap data for the corresponding area, obtained from OSM based on onboard GPS information, is transformed into the ego vehicle’s coordinate system. However, misalignment issues arise due to the low accuracy of the OSM data and biases in the GPS signals, posing a challenge for the fusion of SDMap priors. Following extraction and rasterization, the rasterized SDMap prior often suffers from spatial misalignment, meaning it does not align precisely with the vehicle’s current operational location. This misalignment can be attributed to inaccurate GPS signals or rapid vehicle movement. As a result, the straightforward approach of directly concatenating BEV features with SDMap features becomes ineffective.
Figure 9: The overview of P-MapNet, including BEV feature extractor module, HDMap prior refinement module and BEV feature fusion module.
To address this kind of problem, a multi-head cross-attention module is employed[31], enabling the network to leverage cross-attention to identify the most suitably aligned locations, thus effectively enhancing the BEV features with the SDMap prior in Fig. 9. The SDMap is processed through a convolutional network alongside sine positional embedding, resulting in the generation of SDMap prior tokens. Subsequently, multi-head cross-attention is applied to enrich the BEV queries by incorporating information from the SDMap priors. To improve the continuity and realism of HDMap generation in these scenarios, researchers utilize an adapted pre-trained MAE module[29] to closely approximate the distribution of the HDMap. This approach effectively captures the inherent distribution, enhancing the overall quality of the generated map.
Figure 10: The FlexMap pipeline using GNSS and OSM data.
FlexMap Fusion[66] is composed of three distinct modules as illustrated in Fig. 10: First, the map alignment module aligns the HDmap to the vehicle’s real-time kinematic (RTK)-corrected GNSS trajectory within a projected, local coordinate frame. Second, the map conflation module merges available data from OpenStreetMap (OSM)[8] into a vector map (VM). Lastly, in the georeferencing module, local coordinates are projected into global coordinates to produce a georeferenced VM. The following sections detail the implementation of each FlexMap Fusion module.
In the Map Alignment module, the HDmap is synchronized with the vehicle’s RTK-corrected GNSS trajectory in a projected local coordinate frame. To minimize any shift between these trajectories, the initial GNSS measurement is used as the projection origin. A rigid transformation is employed to account for any remaining rotation or translation discrepancies. To correct residual geometric deviations, typically caused by accumulated errors during SLAM, a piecewise linear rubber-sheet transformation is applied.
Once the Map Alignment module has been completed, the HD and OSM maps are mostly geometrically aligned, with only minor deviations. A preprocessing step is applied to the lane map for refinement. A road network matching algorithm is then used to identify corresponding objects between the vector map (VM) and OSM. After identifying these correspondences, relevant information from OSM is transferred to the lane map. During the map alignment process, the SLAM trajectory and PCM / VM are aligned with the projected GNSS trajectory. To enable georeferencing, the UTM projection applied during Map Alignment is reversed, converting the local coordinates into global ones.
Many existing algorithms construct BEV features using multi-perspective images and employ multi-task heads to delineate road centerlines, boundary lines, pedestrian crossings, and other areas. However, performance tends to decline at the far ends of roads, and challenges arise when the primary subject in the image is occluded. To tackle these issues, It can be seen from Fig. 11 that MapVision[36] incorporates both multi-perspective images and SDMap as inputs. SDMap is utilized as a supplementary element, enhancing the understanding of road topology from a BEV perspective while providing map priors over extended distances.
Figure 11: The overview of MapVision, which uses multi-heads.
The framework of MapVision[36] retains the structure of SMERF[2]. In this setup, lane lines are encoded using sine-cosine position encoding, while category information is represented through one-hot encoding. These features are then concatenated and processed by a transformer encoder to extract meaningful map features. Informed by BEVFormer[18], PV features are projected onto BEV, followed by the fusion of image features with SDMap features within the BEV domain. Moreover, auxiliary foreground segmentation tasks are introduced for both PV and BEV to improve feature extraction capabilities. For the cross-attention mechanism, SDMap features are employed as the key and value, while the feature map derived from multi-perspective images serves as the query. The integration of SDMap priors results in a marked improvement in the quality of HDmap construction. To further enhance the map encoder’s comprehension of the map, a pre-training phase is conducted. This pre-training enables the map encoder to learn the conversion process from the position-encoded SDMap to the feature map output, significantly advancing topological reasoning capabilities.
LGMap[69] introduces a symmetric view transformation that combines forward and backward projections to leverage their complementary advantages. The Lift-Splat-Shoot[19] approach capitalizes on depth distribution to model the uncertainty of each pixel’s depth and dense BEV features from BEVFormer[18], using depth supervision solely during the training phase. This helps mitigate false correlations between 3D and 2D spaces caused by occlusions. By employing sinusoidal embedding, BEVFormer[18] applies cross-attention between the SDMap feature representation and the features from visual inputs at each encoder layer.
Figure 12: The pipeline of LGmap, including symmetric VT module, Lane-TE topology module and Lane-Lane topology module.
In Fig. 12, Hierarchical Temporal Fusion (HTF) fully utilizes the local fusion capabilities of a streaming strategy while also harnessing the long-range fusion strengths of a stacking strategy. This method effectively minimizes memory usage and latency costs compared to the stacking approach. To accommodate various map elements with distinct shape priors, the instance-wise detection decoder is extended with additional segmentation tasks.
Extending the range of BEV representation offers substantial benefits for downstream tasks, including topology reasoning, scene understanding, and planning. This extension provides more comprehensive information and enhances reaction times. The SDMap serves as a lightweight representation of road structure topology, notable for its ease of acquisition and low maintenance costs.
By merging close-range visual data from onboard cameras with beyond line-of-sight (BLOS) environmental priors from SDMap, perceptual capabilities can be expanded to a range of 200 meters. For the BEV feature extraction baseline, LSS[19] is chosen for its lightweight and efficient design, which facilitates easy integration. The backbone of the SDMap Encoder is built upon a VGG architecture.
Figure 13: The pipeline of BLOS-BEV, where the rasterized map is used to generate SDMap features by CNN to participate in feature fusion.
As illustrated in Fig. 13, BLOS-BEV[65] is currently investigating various fusion strategies to combine visual BEV features with SDMap semantics, aiming for optimal representation and performance. The three main approaches being explored are addition, concatenation, and the cross-attention mechanism. Within the BEV Decoder, high and low-resolution fused features are received. The low-resolution features undergo upsampling by a factor of 4 to ensure alignment in height and width with the high-resolution features. Following this, the two sets of features are concatenated along the channel dimension. This is then processed through two convolutional layers and additional upsampling to produce the final BEV segmentation map.
UniHDMap[71] introduces a unified detection framework that integrates information from lanes, pedestrian crossings, and road boundaries. It modifies LaneSegNet[78] for road segmentation, ensuring a unified representation of all road detection elements. To accelerate convergence and improve detection accuracy, it adopts a one-to-many approach by increasing the number of queries fivefold for additional supervision. YOLOv8[79] is utilized for 2D traffic detection, while two MLP-based heads are employed for lane-lane and lane-traffic topology prediction, enhancing the model’s ability to manage complex road scenarios.
Figure 14: The pipeline of UniHDMap.
In Fig. 14, UniHDMap[71] utilizes a ResNet backbone[20] to extract feature maps from images, followed by a view conversion through BEVFormer’s PV-to-BEV encoder module. The transformer-based detection decoders collect BEV features and update queries across layers. During this process, vector-encoded information from standard-definition maps is injected to provide location guidance for map elements. SDMap is converted into vector polyline sequences, which interact with BEV queries via cross-attention after being encoded by a transformer encoder, producing enhanced BEV features for improved map-based detection.
LSM[70] constructs BEV features from multi-view input images using BEVFormer[18]. An SDMap encoder then extracts features from the SDMap elements, which are represented as polylines. As shown in Fig. 15, polylines from the SDMap are encoded into embeddings, followed by cross-attention to fuse the SDMap features with the BEV features.
Additionally, a novel ensemble method is introduced, which decouples topological relationships from road element detection using an MLP. This improves the accuracy of topology prediction, further enhancing model performance in lane detection tasks.
A dynamic positional encoding scheme is proposed for each lane attention layer in the lane decoder, as shown in Fig. 16. This scheme updates the positional encoding by utilizing the location information of lane points derived from the output of the preceding decoder layer. This dynamic update ensures that the positional encoding adapts based on the current state of the lane points, improving the model’s ability to capture and predict the topological structure and spatial relationships of lane elements during decoding.
Figure 15: The pipeline of LSM, including SDMap Encoder and Lane Decoder.
Figure 16: The dynamic positional encoding scheme of LSM.
To enhance performance, an ensemble strategy is applied, starting with the results from the best-performing detection model on the validation set as the base proposals. Other candidate models’ outputs are integrated using a trust-based voting strategy. Depending on the similarity between candidate proposals and base ones, confidence scores are assigned. This approach reduces false negatives and strengthens true positive detection confidence.
RoadPainter[72] processes surrounding images and optionally incorporates SDMap to generate BEV features. These features are used to detect lane centerlines as geometric points in the BEV space and determine their associations, improving the accuracy and reliability of lane topology. To enhance this process, RoadPainter[72] utilizes an SDMap interaction module that augments BEV features with beyond-visual-range data and road shape priors extracted from the SDMap. The SDMap is vectorized and converted into BEV features by filling each grid cell with the corresponding semantic type embedding if occupied.
For extracting lane centerlines, RoadPainter[72] employs a transformer decoder with a hybrid attention layer, which integrates masked cross-attention, deformable cross-attention, and self-attention. Additionally, it uses a topology head based on centerline instance queries and positional embeddings to achieve accurate topological associations.
Figure 17: The pipeline of RoadPainter, which shows three important modules.
To refine the detected centerline points, RoadPainter[72] introduces a points-mask optimization module as illustrated in Fig. 17, consisting of two submodules: points-guided mask generation and points-mask fusion. This method addresses accuracy issues, particularly in areas of high curvature, by leveraging detected centerline points to guide the generation of mask queries. The first stage, mask points sampling, selects a set of points from the generated mask. In the second stage, points fusion combines these sampled mask points with the detected centerline points to produce refined centerline positions. This refinement approach ensures more precise and reliable lane topology reasoning.
EORN[73] centers on utilizing lightweight and scalable SDMap to develop online vectorized HDmap representations. Initially, the integration of rasterized SDMap into various online mapping architectures is explored. A key takeaway from this process is that SDMap encoders are model-agnostic, meaning they can be quickly adapted to new architectures that use BEV encoders. Fig. 18 demonstrates that incorporating SDMap as priors in online mapping tasks significantly accelerates convergence and improves the performance of tasks like online centerline perception.
Figure 18: The overview of ENOR.
To evaluate the utility of SDMap as priors in HD mapping, EORN[73] integrates them into recent online mapping efforts. These tasks fall into two categories: the first focuses on perception, which involves detecting map elements such as lane lines, road boundaries, crosswalks, and centerlines. The second area expands perception to include reasoning, detecting traffic elements like lights and signs, as well as understanding their relationships. In terms of architecture, EORN[73] incorporates SDMap into SOTA HD mapping systems based on MapTR[37]. An image encoder transforms surround-view images into PV features, which are further processed into unified BEV features using a BEV view transform module. A ResNet-18[20] is employed as the SD encoder to extract SDMap features, which are then interpolated and concatenated with the BEV features from the images along the channel dimension, ensuring proper spatial alignment and boosting the mapping performance.
4.1 Alignment Methods